OpenStudy (anonymous):

Find the equation of the line that passes through (5,0) and has a slope of -2/3

OpenStudy (anonymous):

y-y1=m(x-x1) So, y-0=(-2/3)(x-5) =>y=(-2/3)x+ (10/3)

OpenStudy (anonymous):

USE SLOPE INTERCEPT FORM!

OpenStudy (anonymous):

The final answer is in slop intercept form.